OpenStudy (anonymous):

Are people stupid? Why must I do a "double integral" on this problem instead of 2 single ones?

OpenStudy (anonymous):

What I'm trying to find is the area bounded between the curves y=1/x and y=sqrtx inside the domain 1<x<3. The book is telling me to use a double integral, but can't I just take the integral from x=1 to x=3 of y=sqrtx and subtract the integral from x=1 to x=3 of y=1/x to get the area? Doesn't make sense!

OpenStudy (experimentx):

\[ \iint_S dA = A\]

OpenStudy (jhannybean):

you like the integral sign much? :P

OpenStudy (experimentx):

if you do manipulate it carefully, this changes into single integral. of course you can do it via single integral directly.

OpenStudy (amistre64):

as experiment pointed out, a single integral is just a dumbed down version of a double integral: \[\int_{a}^{b}\int_{g(x)}^{f(x)}~~dy~dx\] \[\int_{a}^{b}~f(x)-g(x)~dx\]

OpenStudy (precal):

The more math you studied, the more variations of notation we encounter.